A department store manager has monitored the number of complaints received per week about poor service. The probabilities for numbers of complaints in a week, established by this review, are shown in the table. Number of complaints 0 1 2 3 4 5 Probability 0.15 0.26 0.34 0.06 0.09 0.10 What is the probability of between 2 and 3 (inclusive) complaints received per week? Please specify your answer in decimal terms and round your answer to the nearest hundredth (e.g., enter 12 percent as 0.12).
Answer #1

Given :

The probabilities for numbers of complaints in a week, established by this review, are shown in the table.

x 0 1 2 3 4 5
P(x) 0.15 0.26 0.34 0.06 0.09 0.10

The probability of between 2 and 3 (inclusive) complaints received per week :

P(2 \leq X \leq 3) = P(X=2) + P(X=3)

= 0.34 + 0.06

= 0.40

Therefore the probability of between 2 and 3 (inclusive) complaints received per week is 0.40